Suppose the private marginal cost of producing steel were constant at $45/ton and the marginal damage were increasing MD=2Q, where Q = tons of steel produced.

The social marginal cost of the 100th ton is therefore
Answer #1

Marginal private cost (MPC) = $45 (Constant)

Marginal damage (MD) = 2Q

Marginal social cost = MPC + MD

Marginal social cost = 45 + 2Q

Put Q = 100

Marginal social cost = 45 + 2(100)

Marginal social cost = $245

The social marginal cost of the 100th ton is $245